OpenStudy (anonymous):

In the class elections, Connie received 53 more votes than Carl. In all, 221 votes were cast for the two candidates. How many votes did each receive. One point for the equation, one point for the solution.

OpenStudy (anonymous):

Calling X -> Votes for Connie and Y -> Votes for Carl We know: \[X+Y=221\] \[X-Y=53\] Solving the linear system: \[2X=221+53 \rightarrow X=\frac{274}{2}=137 \] \[Y=221-137=84\] So, Connie received 137 votes and Carl received 84 votes